# Data Generation for Stability Studies of Power Systems with High Penetration of Inverter-Based Resources

### v1

@@ Abstract
Index Terms:
Data generation, High-performance computing, Small-signal stability, Power System
@@ I Introduction
The ongoing energy transition involves the progressive phase-out of fossil-fuel-based Synchronous Generators (SGs), the large-scale integration of renewable Inverter-Based Resources (IBRs), and the increasing electrification of demand. As a consequence, power systems are evolving into power-electronics-dominated systems. In these systems, electromagnetic phenomena become increasingly relevant in shaping the overall dynamic behavior, complementing and in some cases influencing more strongly the traditional electromechanical dynamics that characterize grids dominated by SGs [hatziargyriou2020definition]. These phenomena appear on distinct time scales: electromechanical dynamics evolve relatively slowly, typically over seconds, whereas electromagnetic dynamics occur on the order of microseconds due to the high switching frequencies and fast control loops of power electronic converters [hatziargyriou2020definition]. As a consequence, modern electrical networks are increasingly exposed to fast and potentially poorly damped oscillations that require rapid detection and mitigation. Predicting these events in advance is particularly challenging because of the intermittent and volatile nature of renewable energy sources. To address these new operational challenges, Transmission System Operators (TSOs) are acting both on the monitoring side and on the stability detection and assessment side. On the monitoring side, they are equipping their networks with advanced real-time measurement infrastructures such as Phasor Measurement Units (PMUs) and Wide-Area Monitoring Systems (WAMSs) [CIGRE2023_WAMPAC, SEGUNDOSEVILLA2022107772]. On the assessment side, there is growing interest in Machine Learning (ML)-based surrogate models as alternatives to conventional time-domain simulations [xu2020intelligent]. Conventional simulations provide high accuracy but are computationally expensive. ML-based models, instead, offer the possibility of delivering fast stability assessments suitable for real-time or near–real-time operation [xu2020intelligent].
@@
High-quality training data are essential for the development of ML models for power system stability assessment. Relying solely on historical records from system operation, such as available PMU measurements and corresponding stability labels, is not sufficient to train accurate models, because these data are not fully representative of the operating conditions of interest [xu2020intelligent, thams2019efficient]. Power systems are traditionally operated in a conservative manner, which means that undamped oscillations and instability events are rare. Although the few events that do occur are valuable for understanding the mechanisms that trigger them, they are not sufficient to derive statistically meaningful patterns. ML classifiers, such as those used to predict stability status, require balanced datasets [5128907] and therefore a significant number of unstable operating points (OPs). In addition, high-quality information in the region of the operating space close to the stability boundary is critical, and it has been shown to be the most important factor for obtaining accurate models [giraud2025dataset].
@@
For these reasons, historical data alone cannot provide an adequate training dataset. They must be enriched with synthetic instances generated through simulation. The sampling strategy used to generate such synthetic data should be guided by three main principles: historical relevance, sufficient coverage of the operating space, and discriminative relevance to capture rare events and operating points near the stability margin [bugaje2023generating]. Considering historical relevance, this aspect can be addressed by sampling the operating space according to historical consumption and generation patterns. The study in [gillioz2025large] follows this idea by generating long-term synthetic time series derived from historical national data, producing realistic bus-level power injections for ML applications. To further enhance the representativeness of historically based samples, generative models and other data-driven techniques have been proposed. In [lan2024data], the authors build on historical steady-state data by using Generative Adversarial Network (GAN)-based augmentation and transfer learning to recreate rare operating conditions that are underrepresented in real records. In [xia2024database], historical variability is incorporated by modeling the probability distribution of operating states under uncertain loads and renewable production, generating samples in proportion to their likelihood of occurrence. Although these methods significantly improve realism, they still do not achieve sufficient coverage of the full operating space.
@@
To improve coverage, random sampling techniques such as Monte Carlo sampling and Latin Hypercube Sampling (LHS) are typically employed [duchesne2020recent]. These methods, however, introduce a major challenge: they often produce a large number of infeasible OPs, meaning points for which the power flow (PF) does not converge or for which the solution cannot be used in real operation because it does not satisfy static stability constraints, such as maintaining voltages within acceptable ranges, or because it violates operational limits of generators and transmission lines. The issue of feasibility in sampled OPs is addressed in the studies in [venzke2021efficient, bugaje2023generating]. In [venzke2021efficient], the authors formulate infeasibility certificates based on separating hyperplanes to classify regions of the operating space as infeasible and exclude them from sampling. This removes large portions of the search space and improves scalability, but may also eliminate feasible OPs. The remaining unclassified space is then explored using directed walks to identify the stability margin, followed by sampling from a multivariate distribution fitted to the secure points. Although directed walks initiated from multiple starting points can be parallelized, the overall workflow still follows a fixed sequence of steps: space reduction, directed walks, and sampling in the stability margin. In [bugaje2023split, bugaje2023generating], a power-system adaptation of the split-based GAPSPLIT strategy is proposed. Although modifications are introduced to prevent the algorithm from getting trapped in infeasible regions, the method fundamentally relies on recursively bisecting the input domain. As a result, it remains fully sequential and cannot be parallelized. Moreover, existing methods do not incorporate dimensionality reduction and instead operate in a high-dimensional space (or polytope [venzke2021efficient]), whose dimensionality grows exponentially with system size. Finally, these approaches do not investigate the impact of IBRs, whose presence can significantly jeopardize system stability.
@@
  • A data generation framework tailored to power systems with high IBRs penetration is introduced. The tool is designed to capture how system dynamics evolve under varying penetration levels of grid-forming and grid-following IBRs, as well as under different controller tuning settings.
@@
The remainder of the paper is organized as follows. Section II presents the proposed data generation tool. Section III applies the tool to a case study and evaluates its performance under different parameter settings. Finally, conclusions are drawn in Section IV.
II Data Generation Tool
This section describes the proposed data generation tool, including the input information it requires, how this information is processed, the workflow of the data generation procedure, and the characteristics of the resulting dataset. The main inputs, outputs, and processing steps are summarized in Fig.1. The required inputs include the system information for which the data must be generated, such as generation data, demand profiles, and the grid topology. These inputs can be provided in tabular form (as Excel files), while the network topology can also be supplied through standard RAW files. A detailed description of the input quantities is provided in Section II-A. Based on these inputs, the system’s range of operation is identified and translated into the operating space to be explored. To ensure a scalable exploration of this space, the methodology used to define its dimensions is presented in Section II-B. The operating space is then explored through an iterative data generation process. In each iteration, relevant quantities related to the operating points and controller settings are sampled, and a small-signal stability analysis is performed. Both the sampled variables and the quantities computed during the stability analysis are stored to build the final dataset, which can be exported as a tabular CSV file. The results of each iteration are also processed internally to drive the exploration toward the stability margin, thereby increasing the efficiency and information content of the generated dataset.
@@ II Data Generation Tool
II-A System Setup for Dynamic Studies
@@ II-A1 Range of Operation
II-A2 Equivalents for dynamic studies
To conduct dynamic studies of bulk power systems, it is necessary to simplify the system modeling. In particular, representing each generator individually is impractical; therefore, generator aggregation must be employed. A basic aggregation approach may group generators based on their technology type, such as SGs or IBRs. However, to account for the varying control strategies applied to IBRs (primarily grid-following (GFL) versus grid-forming (GFM)) a different aggregation approach can be implemented, distinguishing generators according to their control mode. To verify the validity of this control-based aggregation approach from a dynamic behavior perspective, an admittance-based frequency scan can be used as a verification method, as shown in Appendix A. Specifically, the conducted frequency response scans prove that the dynamic behavior observed by the grid remains consistent whether all IBRs connected to a bus are modeled individually or aggregated into a single equivalent converter, provided they share the same control strategy. In addition, they reveal that, for identical operating points, the system’s dynamic response differs significantly depending on whether the IBRs are controlled in grid-forming or grid-following mode. This observation highlights the importance of accounting for control strategy when analyzing system stability. It further motivates the generation of a dataset that captures stability assessments across various operating points and different levels of GFL/GFM penetration, while maintaining the same overall operating conditions.
II-B Definition of the Operating Space
@@
Specifically, they include the power injected by each IBR that operates under a control strategy different from the one explicitly sampled (e.g., grid-following ($P_{GFL_{i}}$ $\forall i\in\mathcal{C}$)). These values are determined as the difference between the total IBR power and the portion allocated to IBRs operating in GFM mode, computed on a generator-unit basis.
II-C Space Exploration Workflow
@@
The choice of which dimension to divide when a subregion is split can either be fixed in advance, by specifying the dimensions that should be explored with greater effort, or it can be guided by a sensitivity analysis that identifies the dimension along which the stability classification shows the highest variability. In this second case, a Random Forest is trained using the samples generated up to that point. The model provides an estimate of the most important features involved in the classification process [hastie2009elements], and this information is used to select the dimensions that will be split in the next iteration of the workflow.
II-D Sampling Methods
II-D1 Independent Dimensions Sampling
The sampling of dimension quantities aims to efficiently maximize coverage of the operating space. To this end, the Latin Hypercube Sampling (LHS) technique is applied to all independent dimensions, except for the voltage profile ($\tilde{V}$). The voltage profile must be treated with a different approach compared to the other independent dimensions. Although the goal is to test the system’s behavior under conditions characterized by various voltage profiles, it is important to consider that the voltage at each node cannot be assigned completely at random, as the overall voltage profile of the network is determined by physical constraints and electrical laws governing its operation. Supporting this, it is well known that a network tends to exhibit a natural voltage profile that characterizes its normal operating regime. However, such information may not be available a priori. Therefore, to reconcile random sampling with the need to obtain realistic profiles that reflect the system’s physical behavior, in the absence of detailed knowledge of the conventional voltage profile, the following approach is proposed. A starting bus (e.g., the slack bus) is first selected and assigned a random voltage magnitude within the permissible operating range defined by the grid. The remaining network is then traversed as a graph, assigning to each neighboring bus a voltage magnitude computed as the voltage of the previously visited bus plus a random deviation bounded within a specified range. In the case of meshed networks, where a bus is likely to be reached through multiple paths, its final voltage value is taken as the average of the multiple tentative assignments.
II-D2 Independent Variables Sampling
II-E HPC Application
@@
III Case Study
This section demonstrates the application of the proposed data generation tool to a power system representative of a highly IBR-penetrated grid. The test system is first described, after which datasets are generated both with and without the use of the sensitivity-driven sampling strategy. The resulting datasets are then analyzed to assess their exploration behavior, composition, and effectiveness through the performance of the ML models trained on them. It is worth noting that omitting the sensitivity analysis results in a data generation process that resembles the approach proposed by the authors in earlier work [rossi2022data]. Thus, while the tool presented in this paper is inherently better suited for large systems and high-dimensional operating spaces, the case study is also designed to validate the added value of the sensitivity-based strategy, which represents one of the key novelties of this work.
The simulations are executed in a HPC environment, specifically on the MareNostrum supercomputer at the Barcelona Supercomputing Center. Parallelization and distributed data exploration and generation are managed using PyCOMPSs [tejedor2017pycompss], the Python interface of the COMPSs framework [badia2015comp]. PyCOMPSs relies on a distributed runtime composed of multiple agents that coordinate task scheduling and data management across the computing nodes [lordan2014servicess], enabling transparent and efficient execution of parallel workflows in large-scale HPC environments.
III-A System Setup
The system used to test and validate the data generation tool is the NREL 118-bus system, a modified version of the IEEE 118-bus system characterized by higher installed demand and generation capacity compared to earlier versions. The network is shown in Figure 4. The following subsections present the main features of the system that are relevant for modeling purposes within the data generation framework. For a complete and detailed description, the reader is referred to [pena2017extended], which also provides the full datasets describing the system.
@@ III-A System Setup
III-A1 Generation
TABLE I:
Bus $P_{SG}^{\max}$ $P_{IBR}^{\max}$ $P^{\max}$ $S_{SG,rated}$ $S_{IBR,rated}$ $S_{rated}$ $P_{SG}^{\min}$
$n$ = $P_{SG,nom}$ = $P_{IBR,nom}$ = $P_{SG}^{\max}+P_{IBR}^{max}$ = $\frac{P_{SG,nom}}{\cos\varphi}$ = $\frac{P_{IBR,nom}}{\cos\varphi}$ = $\frac{P_{nom}}{\cos\varphi}$ =20% of $S_{SG,rated}$
$P_{IBR}^{\min}$ $P^{\min}$ $Q^{\min}$ $Q^{\max}$
=20% of $S_{rated}^{IBR}$ =20% of $S_{rated}$ = - $S_{rated}\sin\varphi$ = $S_{rated}\sin\varphi$
@@ III-A1 Generation
The NREL 118-bus system includes 24.6 GW of installed generation capacity and incorporates ten different power generation technologies. In total, there are 327 generators distributed across 54 generation units (i.e., generation buses), as illustrated in Figure 4. Of these, 17 are wind power plants and 35 are photovoltaic solar plants. These renewable facilities represent the IBRs in the system and account for 18.3% of the total installed capacity. The remaining generation capacity is based on SG technology and includes steam turbines (ST) powered by coal, gas, and other fuels; internal combustion engines (ICE) powered by gas; combustion turbines (CT) powered by gas and oil; gas combined-cycle turbines (CC); as well as hydro and biomass generators. The dataset provided in [pena2017extended] specifies, for each generator, its nominal power and the bus to which it is connected. From this information, generators are aggregated by type (SG, IBR, and total) to compute, at each bus, the total nominal and rated power as well as the minimum and maximum active and reactive power injections. The rated power and injection limits are derived from the generator capability curves; for simplicity, all generators are assumed to follow the same curve that sets the active power injection larger than 20% of the rated power, and the reactive power injection with a power factor $\cos\varphi\geq$ 0.95. Table I summarizes the calculation procedure used to obtain each of the above quantities.
III-A2 Demand
@@
The system comprises 91 loads. The dataset provided in [pena2017extended] specifies, for each load, its Load Participation Factor, i.e., the fraction of total system demand it represents.
III-A3 SG and IBR models for small-signal analysis
(a) SG control scheme.
(b) GFOL converter control scheme.
(c) GFOR converter control scheme.
Figure 5: Control schemes
@@ III-A3 SG and IBR models for small-signal analysis
The SG and IBR models considered in this study are those shown in Figure 5. A detailed description of the SG model can be found in [8933043]. The SG model includes the generator, with its electrical circuits represented in the $dq$ reference frame, and a single-mass model capturing the mechanical dynamics. It also incorporates an exciter, a governor, and turbine models.
@@
The GFL converter control scheme includes an inner current control loop, a Phase-Locked Loop (PLL), two outer loops for active and reactive power control, and two droop controls: frequency droop and voltage droop. The GFM converter control, instead, consists of an inner current control loop, outer voltage control loop, and two droop controls: an active power droop (used for synchronization) and a reactive power droop. In both GFL and GFM schemes, each droop controller is preceded by a first-order low-pass filter $G(s)=\frac{1}{\tau s+1}$, where $\tau$ is the filter time constant. The values of $\tau$ for all droop filters ($G_{u}$ and $G_{w}$ in Figures 5(b)5(c)) are considered in this study as independent converter control parameters and are thus treated as control-related independent dimensions.
III-B Analysis of The Generated Datasets

### v2

@@ Abstract
keywords:
Data generation, High-performance computing, Small-signal stability, Power System, High IBRs penetration
journal: Nuclear Physics B
[…]
[label1]organization=Department of Electrical Engineering, CITCEA-UPC,addressline=Avinguda Diagonal, 647, city=Barcelona, postcode=08028, state=Barcelona, country=Spain
[…]
[label2]organization=Barcelona Supercomputing Center, addressline=Plaça d’Eusebi Güell, 1-3, Les Corts, city=Barcelona, postcode=08034, state=Catalonia, country=Spain
@@ 1 Introduction
The ongoing energy transition involves the progressive phase-out of fossil-fuel-based Synchronous Generators (SGs), the large-scale integration of renewable Inverter-Based Resources (IBRs), and the increasing electrification of demand. As a consequence, power systems are evolving into power-electronics-dominated systems. In these systems, electromagnetic phenomena become increasingly relevant in shaping the overall dynamic behavior, complementing and in some cases influencing more strongly the traditional electromechanical dynamics that characterize grids dominated by SGs [1]. These phenomena appear on distinct time scales: electromechanical dynamics evolve relatively slowly, typically over seconds, whereas electromagnetic dynamics occur on the order of microseconds due to the high switching frequencies and fast control loops of power electronic converters [1]. As a consequence, modern electrical networks are increasingly exposed to fast and potentially poorly damped oscillations that require rapid detection and mitigation. Predicting these events in advance is particularly challenging because of the intermittent and volatile nature of renewable energy sources. To address these new operational challenges, Transmission System Operators (TSOs) are acting both on the monitoring side and on the stability detection and assessment side. On the monitoring side, they are equipping their networks with advanced real-time measurement infrastructures such as Phasor Measurement Units (PMUs) and Wide-Area Monitoring Systems (WAMSs) [2, 3]. On the assessment side, there is growing interest in Machine Learning (ML)-based surrogate models as alternatives to conventional time-domain simulations [4]. Conventional simulations provide high accuracy but are computationally expensive. ML-based models, instead, offer the possibility of delivering fast stability assessments suitable for real-time or near–real-time operation [4].
@@
High-quality training data are essential for the development of ML models for power system stability assessment. Relying solely on historical records from system operation, such as available PMU measurements and corresponding stability labels, is not sufficient to train accurate models, because these data are not fully representative of the operating conditions of interest [4, 5]. Power systems are traditionally operated in a conservative manner, which means that undamped oscillations and instability events are rare. Although the few events that do occur are valuable for understanding the mechanisms that trigger them, they are not sufficient to derive statistically meaningful patterns. ML classifiers, such as those used to predict stability status, require balanced datasets [6] and therefore a significant number of unstable operating points (OPs). In addition, high-quality information in the region of the operating space close to the stability boundary is critical, and it has been shown to be the most important factor for obtaining accurate models [7].
@@
For these reasons, historical data alone cannot provide an adequate training dataset. They must be enriched with synthetic instances generated through simulation. The sampling strategy used to generate such synthetic data should be guided by three main principles: historical relevance, sufficient coverage of the operating space, and discriminative relevance to capture rare events and operating points near the stability margin [8]. Considering historical relevance, this aspect can be addressed by sampling the operating space according to historical consumption and generation patterns. The study in [9] follows this idea by generating long-term synthetic time series derived from historical national data, producing realistic bus-level power injections for ML applications. To further enhance the representativeness of historically based samples, generative models and other data-driven techniques have been proposed. In [10], the authors build on historical steady-state data by using Generative Adversarial Network (GAN)-based augmentation and transfer learning to recreate rare operating conditions that are underrepresented in real records. In [11], historical variability is incorporated by modeling the probability distribution of operating states under uncertain loads and renewable production, generating samples in proportion to their likelihood of occurrence. Although these methods significantly improve realism, they still do not achieve sufficient coverage of the full operating space.
@@
To improve coverage, random sampling techniques such as Monte Carlo sampling and Latin Hypercube Sampling (LHS) are typically employed [12]. Using random sampling, however, introduce a major challenge: a large number of infeasible OPs is produced, i.e. points for which the power flow (PF) does not converge or for which the solution cannot be used in real operation because it does not satisfy static stability constraints, such as maintaining voltages within acceptable ranges, or because it violates operational limits of generators and transmission lines. The issue of feasibility in sampled OPs is addressed in the studies in [13, 8]. In [13], the authors formulate infeasibility certificates based on separating hyperplanes to classify regions of the operating space as infeasible and exclude them from sampling. This removes large portions of the search space and improves scalability, but may also eliminate feasible OPs. The remaining unclassified space is then explored using directed walks to identify the stability margin, followed by sampling from a multivariate distribution fitted to the secure points. Although directed walks initiated from multiple starting points can be parallelized, the overall workflow still follows a fixed sequence of steps: space reduction, directed walks, and sampling in the stability margin. In [14, 8], a power-system adaptation of the split-based GAPSPLIT strategy is proposed. Although modifications are introduced to prevent the algorithm from getting trapped in infeasible regions, the method fundamentally relies on recursively bisecting the input domain. As a result, it remains fully sequential and cannot be parallelized. Moreover, existing methods do not incorporate dimensionality reduction and instead operate in a high-dimensional space (or polytope [13]), whose dimensionality grows exponentially with system size. Finally, these approaches do not investigate the impact of IBRs, whose presence can significantly jeopardize system stability.
@@
  • 1.
    A data generation framework tailored to power systems with high IBRs penetration is introduced. The tool is designed to capture how system dynamics evolve under varying penetration levels of grid-forming (GFM) and grid-following (GFL) IBRs, as well as under different controller tuning settings.
  • 2.
@@
The remainder of the paper is organized as follows. Section 2 presents the proposed data generation tool. Section 3 applies the tool to a case study and evaluates its performance under different parameter settings. Finally, conclusions are drawn in Section 4.
2 Data Generation Tool
This section describes the proposed data generation tool, including the input information it requires, how this information is processed, the workflow of the data generation procedure, and the characteristics of the resulting dataset. The main inputs, outputs, and processing steps are summarized in Fig.1. The required inputs include the system information for which the data must be generated, such as generation data, demand profiles, and the grid topology. These inputs can be provided in tabular form (as Excel files), while the network topology can also be supplied through standard RAW files. A detailed description of the input quantities is provided in Section 2.1. Based on these inputs, the system’s range of operation is identified and translated into the operating space to be explored. To ensure a scalable exploration of this space, the methodology used to define its dimensions is presented in Section 2.2. The operating space is then explored through an iterative data generation process. In each iteration, relevant quantities related to the operating points and controller settings are sampled, and a small-signal stability analysis is performed. Both the sampled variables and the quantities computed during the stability analysis are stored to build the final dataset, which can be exported as a tabular CSV file. The results of each iteration are also processed internally to drive the exploration toward the stability margin, thereby increasing the efficiency and information content of the generated dataset.
@@ 2 Data Generation Tool
2.1 System Setup for Dynamic Studies
@@ 2.1.1 Range of Operation
2.1.2 Equivalents for dynamic studies
To conduct dynamic studies of bulk power systems, it is necessary to simplify the system modeling. In particular, representing each generator individually is impractical; therefore, generator aggregation must be employed. A basic aggregation approach may group generators based on their technology type, such as SGs or IBRs. However, to account for the varying control strategies applied to IBRs (primarily GFL versus GFM) a different aggregation approach can be implemented, distinguishing generators according to their control mode. To verify the validity of this control-based aggregation approach from a dynamic behavior perspective, an admittance-based frequency scan can be used as a verification method, as shown in Appendix A. Specifically, the conducted frequency response scans prove that the dynamic behavior observed by the grid remains consistent whether all IBRs connected to a bus are modeled individually or aggregated into a single equivalent converter, provided they share the same control strategy. In addition, they reveal that, for identical operating points, the system’s dynamic response differs significantly depending on whether the IBRs are controlled in GFM or GFL mode. This observation highlights the importance of accounting for control strategy when analyzing system stability. It further motivates the generation of a dataset that captures stability assessments across various operating points and different levels of GFL/GFM penetration, while maintaining the same overall operating conditions.
2.2 Definition of The Operating Space
@@
Specifically, they include the power injected by each IBR that operates under a control strategy different from the one explicitly sampled (e.g., GFL ($P_{GFL_{i}}$ $\forall i\in\mathcal{C}$)). These values are determined as the difference between the total IBR power and the portion allocated to IBRs operating in GFM mode, computed on a generator-unit basis.
2.3 Sampling Methods
2.3.1 Independent Dimensions Sampling
The sampling of dimension quantities aims to efficiently maximize coverage of the operating space. To this end, the LHS technique is applied to all independent dimensions, except for the voltage profile ($\tilde{V}$). The voltage profile must be treated with a different approach compared to the other independent dimensions. Although the goal is to test the system’s behavior under conditions characterized by various voltage profiles, it is important to consider that the voltage at each node cannot be assigned completely at random, as the overall voltage profile of the network is determined by physical constraints and electrical laws governing its operation. Supporting this, it is well known that a network tends to exhibit a natural voltage profile that characterizes its normal operating regime. However, such information may not be available a priori. Therefore, to reconcile random sampling with the need to obtain realistic profiles that reflect the system’s physical behavior, in the absence of detailed knowledge of the conventional voltage profile, the following approach is proposed. A starting bus (e.g., the slack bus) is first selected and assigned a random voltage magnitude within the permissible operating range defined by the grid. The remaining network is then traversed as a graph, assigning to each neighboring bus a voltage magnitude computed as the voltage of the previously visited bus plus a random deviation bounded within a specified range. In the case of meshed networks, where a bus is likely to be reached through multiple paths, its final voltage value is taken as the average of the multiple tentative assignments.
2.3.2 Independent Variables Sampling
2.4 Space Exploration Workflow
@@
The choice of which dimension to divide when a subregion is split can either be fixed in advance, by specifying the dimensions that should be explored with greater effort, or it can be guided by a sensitivity analysis that identifies the dimension along which the stability classification shows the highest variability. In this second case, a Random Forest is trained using the samples generated up to that point. The model provides an estimate of the most important features involved in the classification process [15], and this information is used to select the dimensions that will be split in the next iteration of the workflow.
2.5 HPC Application
@@
3 Case Study
This section demonstrates the application of the proposed data generation tool to a power system representative of a highly IBR-penetrated grid. The test system is first described, after which datasets are generated both with and without the use of the sensitivity-driven sampling strategy. The resulting datasets are then analyzed to assess the process used to generate them, their composition, and their quality through the performance of the ML models trained on them. It is worth noting that omitting the sensitivity analysis results in a data generation process that resembles the approach proposed by the authors in earlier work [16]. Thus, while the tool presented in this paper is inherently better suited for large systems and high-dimensional operating spaces, the case study is also designed to validate the added value of the sensitivity-based strategy, which represents one of the key novelties of this work.
The simulations are executed in a HPC environment, specifically on the MareNostrum supercomputer at the Barcelona Supercomputing Center. Parallelization and distributed data exploration and generation are managed using PyCOMPSs [17], the Python interface of the COMPSs framework [18]. PyCOMPSs relies on a distributed runtime composed of multiple agents that coordinate task scheduling and data management across the computing nodes [19], enabling transparent and efficient execution of parallel workflows in large-scale HPC environments.
3.1 System Setup
The system used to test and validate the data generation tool is the NREL 118-bus system, a modified version of the IEEE 118-bus system characterized by higher installed demand and generation capacity compared to earlier versions. The network is shown in Figure 4. The following subsections present the main features of the system that are relevant for modeling purposes within the data generation framework. For a complete and detailed description, the reader is referred to [20], which also provides the full datasets describing the system.
@@ 3.1 System Setup
Generation
Table 1: Summary of the calculation procedure used to set the generation operational range.
Bus $P_{SG}^{\max}$ $P_{IBR}^{\max}$ $P^{\max}$ $S_{SG,rated}$
$n$ = $P_{SG,nom}$ = $P_{IBR,nom}$ = $P_{SG}^{\max}+P_{IBR}^{max}$ = $\frac{P_{SG,nom}}{\cos\varphi}$
$S_{IBR,rated}$ $S_{rated}$ $P_{SG}^{\min}$ $P_{IBR}^{\min}$ $P^{\min}$
= $\frac{P_{IBR,nom}}{\cos\varphi}$ = $\frac{P_{nom}}{\cos\varphi}$ =20% of $S_{SG,rated}$ =20% of $S_{rated}^{IBR}$ =20% of $S_{rated}$
$Q^{\min}$ $Q^{\max}$
= - $S_{rated}\sin\varphi$ = $S_{rated}\sin\varphi$
@@
The NREL 118-bus system includes 24.6 GW of installed generation capacity and incorporates ten different power generation technologies. In total, there are 327 generators distributed across 54 generation units (i.e., generation buses), as illustrated in Figure 4. Of these, 17 are wind power plants and 35 are photovoltaic solar plants. These renewable facilities represent the IBRs in the system and account for 18.3% of the total installed capacity. The remaining generation capacity is based on SG technology and includes steam turbines powered by coal, gas, and other fuels; internal combustion engines powered by gas; combustion turbines powered by gas and oil; gas combined-cycle turbines; as well as hydro and biomass generators. The dataset provided in [20] specifies, for each generator, its nominal power and the bus to which it is connected. From this information, generators are aggregated by type (SG, IBR, and total) to compute, at each bus, the total nominal and rated power as well as the minimum and maximum active and reactive power injections. The rated power and injection limits are derived from the generator capability curves; for simplicity, all generators are assumed to follow the same curve that sets the active power injection larger than 20% of the rated power, and the reactive power injection with a power factor $\cos\varphi\geq$ 0.95. Table 1 summarizes the calculation procedure used to obtain each of the above quantities.
Demand
@@
The system comprises 91 loads. The dataset provided in [20] specifies, for each load, its Load Participation Factor, i.e., the fraction of total system demand it represents.
SG and IBR models for small-signal analysis
@@
The SG and IBR models considered in this study are those shown in Figure 5-6. A detailed description of the SG model can be found in [21]. The SG model includes the generator, with its electrical circuits represented in the $dq$ reference frame, and a single-mass model capturing the mechanical dynamics. It also incorporates an exciter, a governor, and turbine models.
Figure 5: SG control scheme.
@@
The GFL converter control scheme includes an inner current control loop, a Phase-Locked Loop (PLL), two outer loops for active and reactive power control, and two droop controls: frequency droop and voltage droop. The GFM converter control, instead, consists of an inner current control loop, outer voltage control loop, and two droop controls: an active power droop (used for synchronization) and a reactive power droop. In both GFL and GFM schemes, each droop controller is preceded by a first-order low-pass filter $G(s)=\frac{1}{\tau s+1}$, where $\tau$ is the filter time constant. The values of $\tau$ for all droop filters ($G_{u}$ and $G_{w}$ in Figures 6(a)6(b)) are considered in this study as independent converter control parameters and are thus treated as control-related independent dimensions.
(a) GFL converter control scheme.
(b) GFM converter control scheme.
Figure 6: Converters control schemes.
3.2 Analysis of The Generated Datasets